40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15718     Entry time: Wed Dec 9 12:02:04 2020
Author: gautam 
Type: Update 
Category: LSC 
Subject: POX locking still unsatisfactory 

Continuting the IFO recovery - I am unable to recover similar levels of TRX RIN as I had before. Attachment #1 shows that the TRX RIN is ~4x higher in RMS than TRY RIN (the latter is commensurate with what we had previously). The excess is dominated by some low frequency (~1 Hz) fluctuations. The coherence structure is confusing - why is TRY RIN coherent with IMC transmission at ~2 Hz but not TRX? But anyways, doesn't look like its intensity fluctuations on the incident light (unsurprisingly, since the TRY RIN was okay). I thought it may be because of insufficient low-frequency loop gain - but the loop shape is the same for TRX and TRY. I confirmed that the loop UGF is similar now (red trace in Attachment #2) as it was ~1 month ago (black trace in Attachment #2). Seismometers don't suggest excess motion at 1 Hz. I don't think the modulation depth at 11 MHz is to blame either. As I showed earlier, the spectrum of the error point is comparable now as it was previously.

What am I missing?

Attachment 1: armRIN.pdf  160 kB  | Hide | Hide all
armRIN.pdf
Attachment 2: POX_OLTF.pdf  14 kB  | Hide | Hide all
POX_OLTF.pdf
ELOG V3.1.3-